1) Find the half range Fourier cosine and sine series for

š‘“(š‘¤) = 1 āˆ’ š‘¤; 0 < š‘¤ < 1


1
Expert's answer
2021-06-08T11:53:22-0400

The Fourier series of theĀ half range even functionĀ is given by:


"f(w)=\dfrac{a_0}{2}+\displaystyle\sum_{n=1}^\infin a_n\cos (\dfrac{n\pi w}{1})"

"a_0=\dfrac{2}{1}\displaystyle\int_{0}^{1}(1-w)dw=2[w-\dfrac{w^2}{2}]\begin{matrix}
 1 \\
 0
\end{matrix}=1"

"a_n=\dfrac{2}{1}\displaystyle\int_{0}^{1}(1-w)\cos (\dfrac{n\pi w}{1})dw"

"\int \cos(n\pi w)dw=\dfrac{1}{n\pi}\sin(n\pi w)+C_1"

"\int w\cos(n\pi w)dw"

"=\dfrac{1}{n\pi}w\sin(n\pi w)-\dfrac{1}{n\pi}\int\sin(n\pi w)dw"

"=\dfrac{1}{n\pi}w\sin(n\pi w)+\dfrac{1}{n^2\pi^2}\cos(n\pi w)+C_2"

"a_n=2\displaystyle\int_{0}^{1}(1-w)\cos (n\pi w)dw"

"=2[\dfrac{1}{n\pi}\sin(n\pi w)-\dfrac{1}{n\pi}w\sin(n\pi w)]\begin{matrix}
 1 \\
 0
\end{matrix}"

"-2[\dfrac{1}{n^2\pi^2}\cos(n\pi w))]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{4}{(2k+1)^2\pi^2}"

"b_n=0"


"f(w)=\dfrac{1}{2}+\displaystyle\sum_{k=0}^\infin \dfrac{4}{(2k+1)^2\pi^2}\cos ((2k+1)\pi w)"



The Fourier series of theĀ half range odd functionĀ is given by:


"f(w)=\displaystyle\sum_{n=1}^\infin b_n\sin (\dfrac{n\pi w}{1})"

"b_n=\dfrac{2}{1}\displaystyle\int_{0}^{1}(1-w)\sin (\dfrac{n\pi w}{1})dw"

"\int \sin(n\pi w)dw=-\dfrac{1}{n\pi}\cos(n\pi w)+C_1"

"\int w\sin(n\pi w)dw""=-\dfrac{1}{n\pi}w\cos(n\pi w)+\dfrac{1}{n\pi}\int\cos(n\pi w)dw"

"=-\dfrac{1}{n\pi}w\cos(n\pi w)+\dfrac{1}{n^2\pi^2}\sin(n\pi w)+C_2"

"b_n=2\displaystyle\int_{0}^{1}(1-w)\sin (n\pi w)dw"

"=2[-\dfrac{1}{n\pi}\cos(n\pi w)+\dfrac{1}{n\pi}w\cos(n\pi w)]\begin{matrix}
 1 \\
 0
\end{matrix}"

"-2[\dfrac{1}{n^2\pi^2}\sin(n\pi w))]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{2}{n\pi}"


"f(w)=\displaystyle\sum_{n=1}^\infin\dfrac{2}{n\pi}\sin (n\pi w)"